Bunnell Corporation is a manufacturer that uses job-order costing. On January 1, the companys inventory balances were as fol14. What is the gross margin for the year? Gross margin
Answer #1

Applied overhead = 11.50*41000 = 471500

Actual overhead = 310000+150000 = 460000

Over applied overhead = 471500-460000 = 11500

Adjusted cost of goods sold = 1382600-11500 = 1371100

Gross margin = 3202500-1371100 = 1831400